OpenStudy (anonymous):

What is the constant of variation for the quadratic variation? x 2 3 4 5 6 y 20 45 80 125 180 A. 5 B. 10 C. 15 D. 25

OpenStudy (bloomlocke367):

Do you know what the constant of variation is? Like, what it's asking you?

OpenStudy (anonymous):

not really

OpenStudy (bloomlocke367):

Since k is constant (the same for every point), we can find k when given any point by dividing the y-coordinate by the x-coordinate. For example, if y varies directly as x , and y = 6 when x = 2 , the constant of variation is k = 3 . Thus, the equation describing this direct variation is y = 3x

OpenStudy (igreen):

Wait..it says 'quadratic variation

OpenStudy (igreen):

I believe that's represented by y = kx^2

OpenStudy (bloomlocke367):

yes.

OpenStudy (igreen):

Let's plug in 5 for 'k' and 2 for 'x' and see if we get 20 for 'y'. \(y = kx^2\) \(y = (5)(2^2)\) What's 2^2? @bcimkenz

OpenStudy (anonymous):

4

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yes, and what's 5 * 4?

OpenStudy (anonymous):

20

OpenStudy (igreen):

Yes, so 5 fits as the constant of variation, so 5 is our answer.